Assume that females have pulse rates that are normally distributed with a mean of mu equals 76.0
beats per minute and a standard deviation of sigma equals 12.5
beats per minute. . If 1 adult female is randomly​ selected, find the probability that her pulse rate is less than 79 beats per minute.
The probability is ...
1
Expert's answer
2019-03-22T12:13:41-0400

We are given that females have pulse rates that are normally distributed with a


"\mu=76, \sigma=12.5.""z={{\overline{x}-\mu} \over {\sigma / \sqrt{n}}}""z={{79-76} \over {12.5 / \sqrt{1}}}=0.24"

We have to find


"P(\overline{x}<79)=P(z<0.24)=0.5948."

The probability is


"P(\overline{x}<79)=0.5948"